OpenStudy (anonymous):

A local civic theater has 22 seats in the first row and 21 rows in all. Each successive row contains 3 additional seats. How many seats are in the civic theater?

OpenStudy (anonymous):

21 ∑ k+ 3 = (22+3) + (23 + 3)... k = 22

OpenStudy (anonymous):

so how would you figure out a solid number?